OpenStudy (anonymous):

simplify and express using positive integers?(x^8/y^3)^-1/8

OpenStudy (anonymous):

*positive exponents

OpenStudy (saifoo.khan):

First of all flip them to get rid of the negative ezponent.

OpenStudy (saifoo.khan):

Exponent*

OpenStudy (saifoo.khan):

\[\Huge (\frac{y^3}{x^8})^\frac18\]

OpenStudy (saifoo.khan):

Then, we are going to solve further by multiplying the exponents.

OpenStudy (saifoo.khan):

\[\Huge \frac{y^{3 \times \frac18}}{x^{8 \times \frac 18}}\]

OpenStudy (saifoo.khan):

\[\Huge \frac{x^\frac38}{y}\]ANSWER.

OpenStudy (chaise):

Come on chat please Saifoo if you arn't busy?